OpenStudy (anonymous):

5i/6-2i simplify the expression

OpenStudy (anonymous):

multiply by the conjugate

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

plz show na

OpenStudy (anonymous):

\[\frac{ 5i }{ 6-2i }\left( \frac{ 6+2i }{ 6+2i } \right)\]

OpenStudy (anonymous):

ok is the solution sir

OpenStudy (anonymous):

no thats not the solution yet multiply it out

OpenStudy (anonymous):

ok sir plz show na

OpenStudy (anonymous):

\[\frac{ 30i+10i^2 }{ 36-4i^2 }\]

OpenStudy (anonymous):

that turns into: \[\frac{ 30i+10(-1) }{ 36-4(-1) }\]

OpenStudy (anonymous):

thanks for the solution n help sir

OpenStudy (anonymous):

well what did you get for your final answer

OpenStudy (anonymous):

67/18+30i

OpenStudy (anonymous):

sir m i correct

OpenStudy (anonymous):

\[\frac{ 30i-10 }{ 40 }\]

OpenStudy (anonymous):

rewrite it as: \[\frac{ -10 }{ 40 }+\frac{ 30i }{ 40 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it reduces to: \[-\frac{ 1 }{ 4 }+\frac{ 3i }{ 4 }\]

OpenStudy (anonymous):

^thats your final answer in a+bi form

OpenStudy (anonymous):

thanks a lot i got it

OpenStudy (anonymous):

np

OpenStudy (anonymous):

y